# The convexity of a planar domain via properties of solutions to the modified Helmholtz equation

### v1

@@ 1 Introduction
Let a point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ belongs to a half-plane, say $x_{2}>0$. It is obvious that the distance from $x$ to the $x_{1}$-axis bounding this half-plane — the length of the segment orthogonal to the $x_{1}$-axis is equal to $x_{2}$ — is a harmonic function in this domain. The converse assertion is far from being trivial to prove; it is referred to as one of “three secrets about harmonic functions” in the interesting note [4]. For an arbitrary planar domain $D$ the distance function with respect to $\partial D$ is
@@
where $x$ is a point in $D$ and $|x-y|=\sqrt{(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}}$ is the Euclidean norm in $\mathbb{R}^{2}$. The superharmonicity of this function in $D$ implies that this domain is convex; see Theorem 1.3 below (it was obtained by Armitage and Kuran [3] in 1985). In the present note, it is shown that real-valued solutions of the modified Helmholtz equation (they are labelled panharmonic functions in [5]) also provide a characterization of convexity.
@@
here and below $\nabla=(\partial_{1},\partial_{2})$ is the gradient operator, $\partial_{i}=\partial/\partial x_{i}$. It is known (see [8, p. 437]) that a solution of this problem, say $v(x,\mu)$, satisfies the inequalities
@@
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If the conditions
$v(x,\mu)\neq 0\ \ and\ \ |\nabla v(x,\mu)|\leqslant\mu v(x,\mu),\ \ x\in D,$ (1.4)
are fulfilled for solutions of problem (1.1), (1.2) with all large $\mu>0$, then $D$ is convex.
This theorem is proved in Section 2, and here we illustrate it with an elementary example in which the domain is a disc, and so convex.
Let $I_{0}$ be the modified Bessel function of the first kind (see, for example, [5, p. 111]), then $v(x,\mu)=aI_{0}(\mu|x|)$ with $a\in(0,1)$ solves problem (1.1), (1.2) in the disc $B_{r}(0)=\{y\in\mathbb{R}^{2}:|y|<r\}$ provided $r>0$ is such that $aI_{0}(\mu r)=1$.
It is clear that properties of the function $I_{0}$ (see [2, sect. 9.6]) yield the first condition (1.4) for $v(\cdot,\mu)$, whereas the second condition follows from the inequality $I_{1}(\mu|x|)<I_{0}(\mu|x|)$, because
@@
It is worth mentioning that the original Varadhan’s result is valid for more general elliptic equations, but its proof is rather complicated. In the recent preprint [6], Theorem 1.2 is generalized by adding the source term to equation (1.1); moreover, the rate of convergence of $-\mu^{-1}\log v(x,\mu)$ to $d(x,\partial D)$ is evaluated. The next result has a simple proof which is reproduced here.
@@
Theorem 1.3 is specifically a two-dimensional one because neither $D$ nor $\overline{D}$ need be convex in higher dimensions; an example confirming this is given in [3, Sect. 5].
@@
Let us show that conditions (1.4) imply the superharmonicity of $d(\cdot,\partial D)$ in $D$, which allows us to apply Theorem 1.3, thus yielding the convexity of $D$.
@@
Therefore, if $v(\cdot,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2) with large $\mu>0$, then
@@
provided conditions (1.4) are valid. Hence, $-\mu^{-1}\log v(\cdot,\mu)$ is superharmonic in $D$ for all large $\mu>0$.
@@
because the function on the left-hand side of (2.1) is superharmonic and the first term on the right is harmonic. Letting $\mu\to\infty$ first and then $\rho\to+0$, we arrive at
@@
Since the obtained inequality is valid for all $r>0$ such that $\overline{B_{r}(x)}\subset D$, the distance function $d(\cdot,\partial D)$ is superharmonic in $D$. Then Theorem 1.3 yields the assertion of Theorem 1.1, and this completes the proof.

### v2

@@ 1 Introduction
Let a point $x=(x_{1},x_{2})\in\mathbb{R}^{2}$ belong to a half-plane, say $x_{2}>0$. It is obvious that the distance from $x$ to the $x_{1}$-axis bounding this half-plane — the length of the segment orthogonal to the $x_{1}$-axis — is equal to $x_{2}$, and so is harmonic in this domain. The converse assertion is far from being trivial to prove; it is referred to as one of “three secrets about harmonic functions” in the interesting note [4]. A half-plane is the maximal, as it were, convex domain in $\mathbb{R}^{2}$, whereas our aim is to characterize in a new fashion the convexity of a bounded planar domain.
In an arbitrary planar domain $D$, the distance function with respect to $\partial D$ is
@@
where $x\in D$ and $|x-y|=\sqrt{(x_{1}-y_{1})^{2}+(x_{2}-y_{2})^{2}}$ is the Euclidean norm in $\mathbb{R}^{2}$. It is known that the superharmonicity of this function in $D$ implies that the domain is convex; see Theorem 1.3 below (it was obtained by Armitage and Kuran [3] in 1985). In the present note, we demonstrate that it is also possible to describe the domain’s convexity by means of real-valued solutions of the modified Helmholtz equation (labelled panharmonic functions in [5]); indeed, this is achieved by imposing some conditions on them.
@@
here and below $\nabla=(\partial_{1},\partial_{2})$ is the gradient operator, $\partial_{i}=\partial/\partial x_{i}$.
It has long bean known (see [8, p. 437]) that a solution of this problem, say $v(x,\mu)$, satisfies
@@
Moreover, the left inequality is strict; see [6, Sect. 2.2] for the proof.
@@
Let $D\subset\mathbb{R}^{2}$ be a bounded domain with smooth boundary. If the condition
$|\nabla v(x,\mu)|\leqslant\mu v(x,\mu),\ \ x\in D,$ (1.4)
holds for solutions of problem (1.1), (1.2) with all large $\mu>0$, then $D$ is convex.
This theorem is proved in Section 2, and here we illustrate it with two elementary examples.
In the first one, the domain $D$ is the half-plane $x_{2}>0$, in which case it is natural to require
$v(x,\mu)\to 0\quad\mbox{as}\ x_{2}\to\infty.$
Duffin established that $v(x,\mu)=\mathrm{e}\kern 0.2pt^{-\mu x_{2}}$ solves problem (1.1), (1.2) complemented by the latter condition; see his proof of [5, Theorem 5]. Condition (1.4) is fulfilled for this function for all $\mu>0$, and so Theorem 1.1 implies the obvious fact that $D$ is convex. Notice that
$-\mu^{-1}\log v(x,\mu)=x_{2}=d(x,\partial D)$
in agreement with the assertion of Theorem 1.2 formulated below.
In the second example, the domain $D$ is a disc, and so convex. As usual, $I_{0}$ stands for the modified Bessel function of the first kind (see, for example, [5, p. 111]). If $a\in(0,1)$, then $v(x,\mu)=aI_{0}(\mu|x|)$ solves problem (1.1), (1.2) in the disc $B_{r}(0)=\{y\in\mathbb{R}^{2}:|y|<r\}$ provided $r>0$ is such that $aI_{0}(\mu r)=1$.
It is clear that properties of the function $I_{0}$ (see [2, sect. 9.6]) yield condition (1.4) for $v(\cdot,\mu)$; indeed, it takes the form $I_{1}(\mu|x|)<I_{0}(\mu|x|)$, because
@@
It is worth mentioning that the original Varadhan’s result involves the equation in $\mathbb{R}^{m}$, $m\geqslant 2$, with a general elliptic operator of second order instead of the Laplacian, and the distance is induced by a Riemannian metric derived from the operator’s coefficients. For this reason the Varadhan’s proof is rather complicated. In the recent preprint [6], signed distance functions are considered, thus generalizing $d(\cdot,\partial D)$. For this purpose the source term is added to equation (1.1), which allows the authors to extend Theorem 1.2; moreover, the rate of convergence of $-\mu^{-1}\log v(x,\mu)$ to $d(x,\partial D)$ is evaluated in the preprint.
Furthermore, let us consider $v(x,\mu)$, satisfying equation (1.1) and the Neumann condition
$\partial v/\partial n=\mu\quad\mbox{on}\ \partial D$
($n$ is the exterior normal on $\partial D$), instead of the Dirichlet condition (1.2). It is interesting to find out whether $-\mu^{-1}\log v(x,\mu)$ converges to $d(x,\partial D)$ in $D$ as $\mu\to\infty$ in this case.
Another classical result has a simple proof which is reproduced here.
@@
It should be emphasized that Theorem 1.3 is specifically a two-dimensional result, because neither $D$ nor $\overline{D}$ need be convex in higher dimensions; in [3, Sect. 5], Armitage and Kuran give an example confirming the latter assertion.
@@
It is sufficient to show that conditions (1.4) imply that $d(\cdot,\partial D)$ is superharmonic in $D$; indeed, this allows us to apply Theorem 1.3, thus demonstrating the convexity of $D$.
@@
Therefore, if $v(\cdot,\mu)$ is a solution of the Dirichlet problem (1.1), (1.2) with large $\mu>0$, then it does not vanish, and so
@@
provided condition (1.4) is valid. Hence, $-\mu^{-1}\log v(\cdot,\mu)$ is superharmonic in $D$ for all large $\mu>0$.
@@
Indeed, the function on the left-hand side of (2.1) is superharmonic and the first term on the right is harmonic. Letting $\mu\to\infty$ first and then $\rho\to+0$, we see that the last inequality turns into
@@
Since the obtained inequality is valid for all $r>0$ such that $\overline{B_{r}(x)}\subset D$, the distance function $d(\cdot,\partial D)$ is superharmonic in $D$. Then Theorem 1.3 yields the assertion of Theorem 1.1, thus completing the proof.